OpenStudy (anonymous):

Find the indicated limit, if it exists. limit of f of x as x approaches 1 where f of x equals 1 minus x when x is less than 1, 8 when x equals 1, and x plus 7 when x is greater than 1

OpenStudy (chosenmatt):

what do you think?

OpenStudy (anonymous):

The answers are: 8 0 7 The limit does not exist. I think the answer is 8. Is that correct?

OpenStudy (anonymous):

@freckles

OpenStudy (xapproachesinfinity):

hmmm just take the limit to the right of 1 and also limit to the left of 1

OpenStudy (xapproachesinfinity):

and then compare if they are equal

OpenStudy (chosenmatt):

good job @xapproachesinfinity @sunshine201

OpenStudy (xapproachesinfinity):

so first start with \(\large \lim_{x\to1^{+}}f(x)=\lim_{x\to1^{+}}x+7=?\)

OpenStudy (anonymous):

8?

OpenStudy (xapproachesinfinity):

correct! how about the left

OpenStudy (anonymous):

Is that the one that's 1 - x?

OpenStudy (xapproachesinfinity):

yes x<1 means we are taking number to the left of 1

OpenStudy (anonymous):

so 1 - 8 = -7?

OpenStudy (xapproachesinfinity):

hmmm no we are doing the limit to the left of 1

OpenStudy (xapproachesinfinity):

so you have to evaluate 1 not 8

OpenStudy (xapproachesinfinity):

\(\large \lim_{x\to1^{-}}1-x=?\)

OpenStudy (anonymous):

1 - 1 = 0?

OpenStudy (xapproachesinfinity):

you got it?

OpenStudy (anonymous):

The limit is 0?

OpenStudy (xapproachesinfinity):

yes correct! now is this true \(\large \lim_{x\to1^{+}}f(x)=\lim_{x\to1^{-}}f(x)\)

OpenStudy (xapproachesinfinity):

for the limit to exist the two limits must equal

OpenStudy (xapproachesinfinity):

we are looking for the limit as x goes to 1 so we need to test the limit from the right and the limit from the left of 1 and then compare

OpenStudy (xapproachesinfinity):

limit from the right is 8 limit from the left is 0 what can you say

OpenStudy (xapproachesinfinity):

what do you think

OpenStudy (chosenmatt):

@bohotness